OpenStudy (firechild17):

Solve −2(2x + 5) − 3 = −3(x − 1). 16 −16 −10 −13

OpenStudy (firechild17):

@3mar

likeabossssssss (likeabossssssss):

what do u think it ?

OpenStudy (firechild17):

A?

OpenStudy (brian22):

B

likeabossssssss (likeabossssssss):

no dirct aswer plz

KjSaif (kjsaif):

Step 1: Simplify both sides of the equation. −2(2x+5)−3=−3(x−1) Simplify: (Show steps) −4x−13=−3x+3 Step 2: Add 3x to both sides. −4x−13+3x=−3x+3+3x −x−13=3 Step 3: Add 13 to both sides. −x−13+13=3+13 −x=16 Step 4: Divide both sides by -1. −x/−1=16/−1 x=−16

likeabossssssss (likeabossssssss):

^^

OpenStudy (3mar):

Sorry for being late! It was an emergency call.